ham radio operator often broadcast their signal on a wavelength of 6 meters. what is the frequency of this radio wave
1
Expert's answer
2017-11-30T15:24:07-0500
f = c/l
f = 3*10^8 m/s / 6m = 0.5*10^8 (Hz)